OpenStudy (anonymous):

Use the given degree of confidence and sample data to construct a confidence interval for the population proportion p. A survey of 300 union members in the New York State reveals that 112 favor the republican candidate for governor. Construct the 98% confidence interval for the true population proportion of all New York State union members who favor the Republican candidate.

OpenStudy (amistre64):

\[\frac xn\pm z_{\%}\sqrt{\frac{x(n-x)}{n^3}}\]

OpenStudy (amistre64):

since texts seem to use the pq format these days: p = x/n; q = (n-x)/n which gives us the other setup\[p\pm z_{\%}\sqrt{\frac{pq}{n}}\]